OpenStudy (anonymous):

solve the equation x2+3x-5=0

OpenStudy (anonymous):

you have to use the quadratic formula

OpenStudy (anonymous):

Observe that this is not a perfect square => meaning cross method wouldn't work So if there are 2 ways to solve it now: M1- Completing the Square M2- Graphics calc [APPS -> Plysmlt -> poly root finder] M1: *note that coefficient of x^2 must always be one \[( x ^{2} + 3x ) -5 = 0 \] \[(x + 3/2 )^{2} - (3/2)^{2} - 5 = 0 \] \[(x + 3/2)^{2} -(29/4) \]

OpenStudy (anonymous):

Bascially completing the sq of ax2 + bx + c = 0 a(x2 + bx) + c = 0 a (x + b/2)^2 - a(b/2)^2 + c = 0 *note that if a is not 1, the value that you minus out (b/2)^2 must be a(b/2)^2!